OpenStudy (anonymous):

Use inductive reasoning to find the next term in the sequence 0, 1, 1, 2, 3, 5, 8 … What is the next term?

OpenStudy (queelius):

0+1=1, 1+1-2, 1+2=3, 2+3=5. Do you see the pattern?

OpenStudy (anonymous):

no??